OpenStudy (roadjester):

Determine whether the sequence is increasing, decreasing, or not monotonic. Is the sequence bounded? a_n=n*(-1)^n

OpenStudy (roadjester):

\[a _{n}=n (-1)^n\]

OpenStudy (anonymous):

It is not increasing or decreasing because of the -1

OpenStudy (roadjester):

Sure, but is it bounded?

OpenStudy (anonymous):

\(\{-1,2,-3,4,-5,6,-7,8...\}\)

OpenStudy (kinggeorge):

The absolute value of the sequence is just \(a_n=n\) Is this bounded?

OpenStudy (anonymous):

surely does not look bounded to me

OpenStudy (roadjester):

And I'm taking the absolute value because why?

OpenStudy (roadjester):

I thought I was taking a limit?

OpenStudy (kinggeorge):

Taking the absolute value is just an easier way to see if it's bounded or not. Formally, it's not enough, but it allows you to see why it isn't bounded.

OpenStudy (roadjester):

But I thought the theorem was: \[\lim_{n \rightarrow \infty} \left| a_n \right|=0\] then \[\lim_{n \rightarrow \infty} a_n=0\]

OpenStudy (kinggeorge):

That's correct. Like I said, my explanation is not good enough for a proof. It's merely an intuitive way to see the problem.

OpenStudy (roadjester):

So is it applicable to this situation since the abs is not 0? I mean, since the abs isn't 0, then wouldn't there have to be another way to determine whether it was bounded or not even though intuitively we can tell it isn't?

OpenStudy (kinggeorge):

I might show that the distance between two numbers in your sequence does not approach 0. \[\lim_{n \rightarrow \infty}a_{n+1}-a_n\]

OpenStudy (kinggeorge):

\[=\pm(2n+1)\]